Problem 11-5 Shamrock Mining Company has purchased a tract of mineral land for $1,044,000. It is estimated that this tract wiAs chief accountant for the company, you are to prepare a schedule showing estimated depletion and depreciation costs for eacEstimated depreciation cost Building Machinery (1/2)Machinery (1/2) Year 1st yr. $ 9th Yr. 10th Yr. 11th Yr.Also compute the depreciation and depletion for the first year assuming actual production of 5,800 tons. Nothing occurred dur
Answer #1

1.

Year 1 Estimated depletion cost Depletion 50460 100920 100920 100920 100920 100920 100920 100920 100920 100920 50460

Working as follows:

Year 1 Estimated depletion cost Depletion =(1044000-34800)*(6960/139200) =(1044000-34800)*(13920/139200) =(1044000-34800)*(13